OpenStudy (anonymous):

write the slope-intercept equation for the line that passes through (-4,-8) and (-3,-3) please show work

OpenStudy (anonymous):

find slope of line which is y2-y1/x2-x1. -3-(-8)/-3-(-4) <__double negative cancel to become positvie 5/1=5 y=5x+b if you need to find the y intercept you take one of your points and plug it into slope intercept form equation -8=5(-4)+b -8=-20+b b=12 hope that helps :)

OpenStudy (anonymous):

i messed up it should have been (-4,-8) and (-3,-2)

OpenStudy (anonymous):

okay well its easy just plug your numbers into what i showed u -2+8/-3+4=6/1=6 y=6x+b -8=6(-4)+b -8=-24+b b=16

OpenStudy (anonymous):

thank you. everything is starting to make more sense now

OpenStudy (anonymous):

yeah np glad i could help